OpenStudy (anonymous):

How to write A(x) = 2|x+5| - 3|x-1| as piecewise defined?

OpenStudy (anonymous):

\[2|x+5|=2x+10\] if \(x>-5\) and \(-2x-10\) is \(x<-5\) is a start

OpenStudy (anonymous):

you have to divide up in to the intervals \((-\infty, -5), (-5,1,), (1,5), (5,\infty)\) and check what your function is on each interval